Determine weight loss needed to reach a desired body fat percentage for the following person. A runner weighs 145 pounds and 18% is fat mass and 82% is lean mass. She wants to decrease to 16% fat mass. How much weight will she need to lose to achieve this goal, assuming that all weight lost comes from fat?

Formula: Body weight = Fat Free Mass/(1- desired body fat percentage (decimal))

Weight loss __________________

Using Cunningham’s what was her RMR before weight loss?

What is the RMR now at 16% fat mass.

Determine your weight. Pretend that you want to gain 10 pounds. Develop a plan on how you can achieve this goal.

1) What would be an advisable weight gain rate per week?

2) How many calories do you currently need (before accounting for wanted weight gain)? Guess on your fat free mass if you do not know it already.

Use the Cunningham equation. 500 + (22 X FFM) = ___________

Multiply with a PA factor. _________

Total Caloric need _____________

3) Since you want to gain weight how many more calories would you want to consume per day by the amount per week determined in question one.

Increase caloric intake by ________calories/day.

New total caloric daily need ____________calories/day
Answer #1

1.Weight of the patient is 145 pounds

Fat mass is 18 % = 0.18

Required fat mass is 16% = 0.16

To reduce the fat mass how much weight has to be deducted

= 145 × 0.18 = 26pounds is fat

this meas her lean body mass is 145 - 26 = 119 pounds

fat percentage to be reduced to 16 % = 0.16

substract it from 1 = 1- 0.16 = 0.84

Then divide the lean body mass from 0.84

= 119/ 0.84 = 141 .6 ~ 142 pounds is needed to achieve target fat %

hence , patient need lose 3pounds ( 145 - 142 ) to achieve target fat %of 16%.

2. RMR is resting metabolic rate before weight loss = 500 + fat free mass × 22 ( Cunningham equation)

fat free mas is lean body mass

= 500 + 22× 119

=3118 cal

RMR with 16% fat , weight 145pounds

fat % = 0.16

145× 0.16 = 23.2 ~23 pound is fat

And lean body mass = 145-23 = 122pounds

Now , RMR = 500+ 22× fat free mass

= 500 + 22× 122

= 3184 cals